# Error Signals for Overcoming the Laser Power Limits of Gravitational-Wave Detection

### v1

@@ Abstract
A major barrier to improving the quantum-limited sensitivity of gravitational-wave observatories are the thermal distortions of the test masses which arise at megawatt laser power. Recent advances in a new form of higher-order wavefront correction, in which corrective heating profiles are applied to the test mass surfaces near their edges, have the potential to enable a tenfold reduction of the quantum noise floor of future detectors. However, realizing high levels of quantum noise reduction in practice hinges on identifying measurable error signals to finely control each wavefront actuator, in order to suppress wavefront errors to few-nanometer precision across the full mirror apertures. No direct source of such an error signal exists in LIGO today. We demonstrate that thermally imaging the surface of each test mass can provide these critical error signals. We show that the surface temperature profiles obtained from thermal imaging can be uniquely mapped to a finite element model of the mirror whose complete thermal state is identified, enabling full-aperture wavefront reconstruction and direct error signals for real-time precision wavefront control. This new sensing capability can enable up to a 34% strain sensitivity improvement in LIGO A+ at 95% confidence, increasing the sky-averaged detection range for binary neutron star mergers by 11 Mpc, and will be integral to a next-generation 40-km gravitational-wave observatory in the U.S., Cosmic Explorer.
I Introduction
@@
The achievable arm power and squeezing are limited, in practice, by thermal distortions of the test masses [20]. In recent work [21], we reported the potential of a new class wavefront actuator, known as the FROnt Surface Type Irradiator (FROSTI), to enable next-generation quantum noise targets when used to augment LIGO’s existing thermal compensation capabilities [22]. However, realizing the potential sensitivity improvements of Ref. [21] relies on there also existing error signals to precisely identify the optimal power settings of the wavefront actuators, so as to minimize root mean square (RMS) wavefront errors across the full aperture of the mirrors. No direct error signals containing full-aperture wavefront information exist in LIGO today. LIGO’s existing Hartmann wavefront sensors [22], for example, are fundamentally limited in their field of view to only the central region of the test masses. Similarly, Virgo employs phase cameras as wavefront sensors to measure the spatial structure of the laser fields, but these systems also suffer from a limited field of view due to the limited beam size [23].
@@
In this letter, we present a novel sensing technique that can provide these critical error signals using thermal imaging cameras. We describe this sensing technique in §II and demonstrate its feasibility in §III, assuming realistic sensor resolution limits. We then show in §IV that the improved capability provided by this technique to infer, and thus to more finely correct, thermally-driven wavefront errors in gravitational-wave detectors will significantly improve their quantum-limited strain sensitivity, which we find can be improved up to 34% in LIGO A+. Finally, we discuss the enabling impact of this work on future gravitational-wave detectors, beyond A+, in §V.
II Thermal State Inference
@@
We aim to uniquely and accurately infer the full thermal state of each test mass directly from the temperature profile of its front (reflective-coated) surface. This surface temperature map is measured in situ using a thermal imaging camera, located outside the vacuum system, and matched to a finite element analysis (FEA) thermal model of the test mass. As illustrated in Fig. 1, the thermal steady state is fully characterized by five parameters: the power absorbed from the 1064-nm laser beam, $P_{\rm S}$; the beam’s location on the test mass, $(x_{0},y_{0})$, which in general can be offset from the center; and the absorbed powers delivered by two different wavefront actuators. The actuators include LIGO’s existing ring heater [24], which projects a heating power of $P_{\rm R}$ onto the barrel of the test mass to correct defocus (radius of curvature) error, and the recently proposed FROnt Surface Type Irradiator (FROSTI) [21], which projects an annular heating pattern of power $P_{\rm F}$ onto the front surface of the test mass, near the edge, to correct higher-order aberrations. These parameters are summarized in Table 1.
@@
The functions $\hat{T}_{\rm R}(x,y)$ and $\hat{T}_{\rm F}(x,y)$ respectively represent the surface temperature maps due to the ring heater and FROSTI actuators delivering unit power. Similarly, the function $\hat{T}_{{\rm S}}(x,y\,|\,x_{0},y_{0})$ represents the surface temperature map due to unit power absorbed from the incident laser beam, whose centroid is located at position $(x_{0},y_{0})$ relative to the center of the test mass. We refer to $\hat{T}_{\rm S}$, $\hat{T}_{\rm R}$, and $\hat{T}_{\rm F}$ as unit temperature maps or “unit maps.” The linear addition of these units maps assumed in Eq. 1 is confirmed by FEA modeling to be valid at the power levels of interest ($P_{\rm S}\sim 1\;{\rm W}$ and $P_{\rm R}\sim P_{\rm F}\sim 10\;{\rm W}$).
@@
evaluated over the entire surface of the test mass. The spatial overlap integral of the measured temperature map (Eq. 1) with any unit map, $\hat{T}_{\rm\alpha}$, can thus be expressed as
@@
Figure 1: Test mass thermal state inference from front-surface thermal imaging. It is used to infer the parameters in Table 1.
@@
Table 1: Parameters defining the thermal steady state of one test mass. Each parameter is described further in §II.
Evaluating Eq. 3 for all three unit maps yields the matrix equation
@@
An FEA thermal model of the test mass is used to calculate each unit map, from which the $3\times 3$ matrix of overlap coefficients in Eq. 5 can be numerically calculated.
Although the beam heating unit map in Eq. 5, $\hat{T}_{{\rm S}}$, depends on the beam’s unknown position on the test mass, the absorbed powers and the beam position can be jointly inferred via an iterative procedure. First, we solve Eq. 5 assuming that the beam is perfectly centered, using the unit map $\hat{T}_{{\rm S}}(x,y\,|\,0,0)$. Second, using these initial estimates of the absorbed powers, we subtract the ring heater and FROSTI contributions from the measured temperature map to obtain a residual map representing the beam heating. We estimate the true position of the beam’s centroid as the location of the global maximum in this residual temperature map. Third, we solve Eq. 5 again using the beam heating unit map for the estimated centroid, $\hat{T}_{{\rm S}}(x,y\,|\,x_{0},y_{0})$. Thus, we obtain an estimate of the actual beam position on the test mass, $(x_{0},y_{0})$, from step 2 and then use this to refine our estimate of the absorbed powers, $P_{\rm S}$, $P_{\rm R}$, and $P_{\rm F}$, in step 3. With knowledge of these five parameters, we can use an FEA thermal model of the test mass with their specific values to generate full-aperture mirror phase maps of the actual wavefront errors in reflection and transmission. These phase maps provide a direct error signal for optimizing the ring heater and FROSTI power settings for each test mass.
III Model Performance
@@
Figure 2: Performance of the thermal state inference algorithm. Left: Error of the inferred radial position of the laser beam, as a function of the spatial resolution of the thermal image. Right: Relative errors of the inferred powers, as a function of the temperature resolution of the thermal image.
@@
Fig. 2 shows the model’s ability to accurately recover the true thermal state of the test mass, as the spatial and temperature resolutions of the thermal imaging camera are varied. The left panel shows the average error on the inferred radial position of the laser beam as a function of the spatial resolution (pixelation scale) of the image. The right panel shows the average relative errors on the inferred power levels of the contributing thermal sources as a function of the temperature resolution of the image. We find that a commercial thermal imaging camera, with a $600\times 600$ pixel array and a temperature resolution of 10 mK (indicated by the dashed vertical lines in Fig. 2), paired with an appropriate set of image relay lenses can accurately recover the true beam position to better than 0.5 mm and the true power levels to better than 0.1% for all thermal sources. The feasibility of similar thermal imaging systems has already been demonstrated in Virgo to monitor for point absorbers on the test masses [26].
IV Astrophysical Impact
@@
Given a true set of wavefront errors due to laser beam heating, we use Monte Carlo to simulate our practical inability to perfectly resolve those wavefront errors, which thus limits the precision to which we can identify the optimal power settings of the wavefront actuators and the optimal beam positions. In each trial, we randomly sample the laser beam position, the ring heater power setting, and the FROSTI power setting on each of the four test masses. These parameters are drawn from independent Gaussian distributions centered on their optimal values, corresponding to perfectly centered beams and the wavefront actuator powers which minimize the full-aperture RMS wavefront errors. The standard deviation of each parameter is chosen to match the level of inferential uncertainty in practically determining its optimal value.
For the case with thermal imaging of the test masses, we adopt $3\sigma$ distribution widths of 0.5 mm for the beam positional uncertainties and 0.1% for the ring heater and FROSTI power uncertainties, based on the findings in §III. For the case without thermal imaging, we assume the same inference is performed using only signals available in LIGO today. Accordingly, we adopt a $3\sigma$ distribution width of 1 mm for the beam positional uncertainties, based on the current best knowledge provided by infrared cameras [29]. The laser power absorbed by each test mass can be estimated as the product of the arm power, with an average calibration uncertainty of 3.1% [30], and the absorptivity of the reflective coating, with an average measurement uncertainty of 31.4% [31]. This yields a total relative uncertainty of 35.5% in the power absorbed by each test mass, which we adopt as the $3\sigma$ distribution width of the ring heater and FROSTI powers.
Fig. 3 shows the projected impact of the new wavefront sensing capabilities on the key performance metrics that determine LIGO’s quantum-limited sensitivity. The two cases, with (orange) and without (blue) thermal imaging of the test masses, are each obtained from roughly 1000 Monte Carlo trials assuming the inferential uncertainties described above. We find that 1000 trials are adequate to fully sample the distributions, based on convergence testing of their 95th percentiles. The left panel shows the probability distributions of laser power in the detector’s arm cavities for a fixed input power of 125 W. The middle panel show the probability distributions of observed squeezing at 1 kHz for a fixed injected squeezing level of 9 dB, an effective level chosen to account for all attenuation losses and phase noise from the A+ squeezer system. The right panel shows the probability distributions of total resulting strain noise at 1 kHz. There exist cases in which trade-offs between optimizing the arm power (via the power recycling cavity coupling) and squeezing (via the signal recycling cavity coupling) lead to slightly better performance in one of these metrics, individually, without thermal imaging. However, the overall quantum noise performance is consistently improved when thermal imaging is employed, as shown in the rightmost panel. We find that the higher-precision inference enabled by the thermal imaging results in a dramatic narrowing of the range of possible outcomes, significantly improving the worst possible sensitivity outcome at 95% confidence.
Figure 4: Comparison of the worst-case strain sensitivities of a LIGO A+ detector, at 95% confidence, with and without thermal-imaging-based inference. These two cases correspond to the two cases denoted by the dashed vertical lines in Fig. 3.
This large difference in worst-case scenarios translates directly into a statistical improvement in the expected strain sensitivity of the detector, which is shown in Fig. 4 for the worst observed outcomes in 95% of trials. That is, with 95% confidence, the detector will achieve at least the strain sensitivity indicated by these curves. Each case assumes the nominal A+ input power of 125 W and an effective injected level of 9 dB of frequency-dependent squeezing, as well as the nominal A+ thermal and technical noise contributions which are combined with the quantum noise curves calculated using Finesse. We find that thermal imaging of the test masses can enable up to a 34% greater reduction of the noise floor in LIGO A+, with maximum impact at the highest frequencies, corresponding to an increase of 11 Mpc in the sky-averaged detection range for binary neutron star mergers.
Finally, Fig. 5 shows the detector’s projected performance as the laser power is raised beyond the nominal A+ arm power target of 750 kW, towards the higher power levels required for LIGO $\rm A^{\#}$ and Cosmic Explorer. Each panel contains two shaded regions representing the range of outcomes spanned by the best 95% of trials in each of the two cases, with (orange) and without (blue) thermal imaging of the test masses. The left panel compares the possible ranges of input power required to achieve a given arm power, at 95% confidence. The middle and right panels compare the ranges of possible squeezing and strain sensitivity outcomes, for a given arm power level, at 95% confidence. Without the new inferential capabilities provided by thermal imaging, the probability of significant squeezing and power degradations grows rapidly above 750 kW, which is exacerbated by higher-order spatial modes becoming co-resonant in the recycling cavities with the changing thermal state of the ITMs. With the new sensing capabilities, the quantum noise performance is limited only by the intrinsic capability of the wavefront actuators to match the spatial profiles of the distortions.
V Discussion
Our analysis shows that real-time precision inference of the thermal state of the optics is vital to achieving greater quantum-limited sensitivity in gravitational-wave detectors. We have demonstrated that front-surface thermal imaging offers a practical solution. The requirements for full-aperture wavefront correction are most stringent for the ITMs, due to wavefront distortions shifting the higher-order mode resonances of the signal recycling cavity [20], and the sensing requirements identified in this letter can satisfy them. Our assumption of a thermal steady-state FEA model is a realistic assumption for future detectors. The LIGO A+ upgrades include an offline 10.6-$\mu\rm m$ laser heating system for each ITM, known as the Central HEater for Transient Attenuation (CHETA) [32], designed to maintain the thermal steady state of the optic when the main laser beam is not present due to lock loss. We anticipate that such an offline heating system will be extended to all four test masses by LIGO $\rm A^{\#}$ and will be incorporated into the baseline Cosmic Explorer design. The impact of this work will thus extend far beyond A+, as it also delivers a key enabling sensing capability for next-generation gravitational-wave observatories.
Acknowledgements.

### v2

# Error signals for overcoming the laser power limits of gravitational-wave detectors

@@ Abstract
A major barrier to improving the quantum-limited sensitivity of gravitational-wave observatories is the thermal distortions of the test masses which arise at megawatt laser power. Recent advances in a new form of higher-order wavefront correction, in which corrective heating profiles are applied to the test mass surfaces near their edges, together with other planned instrumental upgrades, have the potential to enable a tenfold reduction of the quantum noise floor of future detectors. However, realizing high levels of quantum noise reduction in practice hinges on identifying measurable error signals to finely control each wavefront actuator, in order to suppress wavefront errors to a few-nanometer precision across the full mirror apertures. No direct source of such an error signal exists in LIGO today. We demonstrate that thermally imaging the surface of each test mass can provide these critical error signals. We show that the surface temperature profiles obtained from thermal imaging can be uniquely mapped to a finite element model of the mirror whose complete thermal state is identified, enabling full-aperture wavefront reconstruction and direct error signals for real-time precision wavefront control. This new sensing capability can enable up to a 34% strain sensitivity improvement in LIGO A+ at 95% confidence, increasing the sky-averaged detection range for binary neutron star mergers by 11 Mpc, and will be integral to a next-generation 40-km gravitational-wave observatory in the U.S., Cosmic Explorer.
keywords:
gravitational-wave detection, thermal compensation, adaptive optics
articletype: Paper
1 Introduction
@@
The achievable arm power and squeezing are limited, in practice, by thermal distortions of the test masses [20]. In recent work [21], we reported the potential of a new class wavefront actuator, known as the FROnt Surface Type Irradiator (FROSTI), to enable next-generation quantum noise targets when used to augment LIGO’s existing thermal compensation capabilities [22]. However, realizing the potential sensitivity improvements of Ref. [21] relies on there also existing error signals to precisely identify the optimal power settings of the wavefront actuators, so as to minimize root mean square (RMS) wavefront errors across the full aperture of the mirrors. No direct error signals containing full-aperture wavefront information exist in LIGO today. LIGO’s existing Hartmann wavefront sensors [22], for example, are fundamentally limited in their field of view by the size of the probe beam. For optical systems with large optics, such as gravitational-wave detectors, this restriction constrains the field of view to only the central region of the optics. Similarly, Virgo employs phase cameras as wavefront sensors to measure the spatial structure of the laser fields, but these systems also suffer from a limited field of view due to the limited beam size [23].
@@
In this paper, we present a novel sensing technique that can provide these critical error signals using thermal imaging cameras for optics in a thermal steady state. This condition will be ensured in future LIGO operation, as the LIGO A+ upgrades include an offline 10.6-$\mu\rm m$ laser heating system for each input test mass (ITM), known as the Central HEater for Transient Attenuation (CHETA) [24], designed to maintain the thermal steady state of the optic when the main laser beam is not present due to lock loss. Furthermore, our sensing technique can be used to match the size and location of the CHETA beam to the interferometer’s main laser beam, as the algorithm runs on millisecond timescales, which is much faster than the typical thermal transient timescale of minutes to hours. We describe this sensing technique in §2 and demonstrate its feasibility in §3, assuming realistic sensor resolution limits. We then show in §4 that the improved capability provided by this technique to infer, and thus to more finely correct, thermally-driven wavefront errors in gravitational-wave detectors will significantly improve their quantum-limited strain sensitivity, which we find can be improved up to 34% in LIGO A+. Finally, we discuss the enabling impact of this work on future gravitational-wave detectors, beyond A+, in §5.
2 Thermal State Inference
@@
We aim to uniquely and accurately infer the full thermal state of each test mass directly from the temperature profile of its front (reflective-coated) surface. This surface temperature map is measured in situ using a thermal imaging camera, located outside the vacuum system, and matched to a finite element analysis (FEA) thermal model of the test mass. The calibrated FEA model of the test mass can then be used to produce mirror phase maps representing the thermally induced wavefront errors arising from both thermoelastic surface deformation and thermorefractive substrate lensing, as uniquely determined by its thermal state. As illustrated in figure 1, the thermal steady state is fully characterized by five parameters: the power absorbed from the 1064-nm laser beam, $P_{\rm S}$; the beam’s location on the test mass, $(x_{0},y_{0})$, which in general can be offset from the center; and the absorbed powers delivered by two different wavefront actuators. The actuators include LIGO’s existing ring heater [25], which projects a heating power of $P_{\rm R}$ onto the barrel of the test mass to correct defocus (radius of curvature) error, and the recently proposed FROnt Surface Type Irradiator (FROSTI) [21], which projects an annular heating pattern of power $P_{\rm F}$ onto the front surface of the test mass, near the edge, to correct higher-order aberrations. These parameters are summarized in table 1.
@@
The functions $\hat{T}_{\rm R}(x,y)$ and $\hat{T}_{\rm F}(x,y)$ respectively represent the surface temperature maps due to the ring heater and FROSTI actuators delivering unit power. Similarly, the function $\hat{T}_{{\rm S}}(x,y\,|\,x_{0},y_{0})$ represents the surface temperature map due to unit power absorbed from the incident laser beam, whose centroid is located at position $(x_{0},y_{0})$ relative to the center of the test mass. We refer to $\hat{T}_{\rm S}$, $\hat{T}_{\rm R}$, and $\hat{T}_{\rm F}$ as unit temperature maps or “unit maps.” The linear addition of these units maps assumed in equation 1 is confirmed by FEA modeling to be valid at the power levels of interest ($P_{\rm S}\sim 1\;{\rm W}$ and $P_{\rm R}\sim P_{\rm F}\sim 10\;{\rm W}$).
@@
evaluated over the entire surface of the test mass. The spatial overlap integral of the measured temperature map (equation 1) with any unit map, $\hat{T}_{\rm\alpha}$, can thus be expressed as
@@
Figure 1: Test mass thermal state inference from front-surface thermal imaging. This procedure is used to infer the parameters in table 1.
@@
Table 1: Parameters defining the thermal steady state of one test mass. Each parameter is described further in §2.
Evaluating equation 3 for all three unit maps yields the matrix equation
@@
An FEA thermal model of the test mass is used to calculate each unit map, from which the $3\times 3$ matrix of overlap coefficients in equation 5 can be numerically calculated.
Although the beam heating unit map in equation 5, $\hat{T}_{{\rm S}}$, depends on the beam’s unknown position on the test mass, the absorbed powers and the beam position can be jointly inferred via an iterative procedure. First, we solve equation 5 assuming that the beam is perfectly centered, using the unit map $\hat{T}_{{\rm S}}(x,y\,|\,0,0)$. Second, using these initial estimates of the absorbed powers, we subtract the ring heater and FROSTI contributions from the measured temperature map to obtain a residual map representing the beam heating. We estimate the true position of the beam’s centroid as the location of the global maximum in this residual temperature map. Third, we solve equation 5 again using the beam heating unit map for the estimated centroid, $\hat{T}_{{\rm S}}(x,y\,|\,x_{0},y_{0})$. Thus, we obtain an estimate of the actual beam position on the test mass, $(x_{0},y_{0})$, from step 2 and then use this to refine our estimate of the absorbed powers, $P_{\rm S}$, $P_{\rm R}$, and $P_{\rm F}$, in step 3.
With knowledge of the three absorbed powers and the laser beam position, direct error signals for the two wavefront actuators can be constructed from the difference between the inferred actuator powers and their optimal values for the inferred level of absorbed laser power ($P_{\rm R,0}$ and $P_{\rm F,0}$ for the ring heater and FROSTI, respectively):
$\varepsilon_{\rm R}$ $=P_{\rm R}-P_{\rm R,0}$ (6)
$\varepsilon_{\rm F}$ $=P_{\rm F}-P_{\rm F,0}$ (7)
Since the optimal actuator powers scale linearly with the absorbed laser power (i.e., the ratios of optimal actuator powers to self-heating power are invariant), these error signals optimize the ring heater and FROSTI power settings for each test mass with minimal computational cost, without requiring running expensive FEA models on the fly. Thus, minimizing these error signals minimizes the actual wavefront errors in reflection and transmission across the full mirror aperture.
3 Model Performance
@@
Figure 2: Performance of the thermal state inference algorithm. Left: Error of the inferred radial position of the laser beam, as a function of the spatial resolution of the thermal image. Middle: Relative errors of the inferred powers, as a function of the temperature resolution of the thermal image. Right: Relative power errors versus the fractional higher-order mode power, shown for the first-order HG1,0 mode (solid lines) and second-order LG1,0 mode (dashed lines), which arise in laser cavities from small misalignment and mode-mismatch, respectively [27].
@@
Figure 2 shows the model’s ability to accurately recover the true thermal state of the test mass, as the spatial and temperature resolutions of the thermal imaging camera are varied. The left panel shows the average error on the inferred radial position of the laser beam as a function of the spatial resolution (pixelation scale) of the image. The middle panel shows the average relative errors on the inferred power levels of the contributing thermal sources as a function of the temperature resolution of the image. We find that a commercial thermal imaging camera, with a $600\times 600$ pixel array and a temperature resolution of 20 mK (indicated by the dashed vertical lines in figure 2), paired with an appropriate set of image relay lenses can accurately recover the true beam position to better than 0.5 mm and the true power levels to better than 0.1% for all thermal sources. The feasibility of similar thermal imaging systems has already been demonstrated in Virgo to monitor for the presence of point absorbers on the test masses [28].
The thermal state inference algorithm assumes that the primary heating beam is a pure fundamental Gaussian mode. The presence of higher-order spatial modes can introduce systematic error that degrades the predictive accuracy. In particular, the first-order Hermite-Gaussian mode (HG1,0) and second-order Laguerre-Gaussian mode (LG1,0) commonly arise in laser cavity experiments due to small misalignment and mode-mismatch, respectively [27]. The right panel of figure 2 shows the effect of such higher-order mode content on the inferred power error. To achieve an inferential accuracy of 0.1%, we find that the fraction of power in higher-order modes must not exceed approximately 1700 ppm, which is readily achievable with modern precision sensing and control techniques [29, 30].
With thermal imaging of the test masses, we estimate that the $3\sigma$ inferential uncertainties of the beam position and the optimal ring heater and FROSTI power settings (given a true level of laser beam heating) reduce to approximately 0.5 mm and 0.1%, respectively, based on the findings in figure 2. This is in comparison to the scenario without thermal imaging, in which we assume the same inference is performed using only signals available in LIGO today. In that case, we estimate a $3\sigma$ inferential uncertainty of 1 mm on the beam position, based on the current best knowledge provided by infrared cameras [31]. The laser power absorbed by each test mass can be estimated as the product of the arm power, with average calibration uncertainty of 3.1% [32], and the absorptivity of the reflective coating, with an average measurement uncertainty of 31.4% [33]. This yields an total relative uncertainty of 35.5% in the power absorbed by each test mass, which we estimate as the $3\sigma$ inferential uncertainty of the optimal ring heater and FROSTI power settings.
Figure 3: Residual wavefront error phase maps for the ITM surface, ITM substrate, and ETM surface (left to right). The top and bottom rows show results with and without thermal-imaging-based sensing, respectively, at their $3\sigma$ maximum power and position errors. The RMS error for each map is indicated. The FROSTI ETM profile includes intentional edge actuation to shift the resonance frequency of a higher-order mode relative to the fundamental mode (see Ref. [26]), but this feature is removed here to better visualize the residual error in the central region.
Figure 3 shows an example of the residual wavefront errors that could arise under these levels of inferential uncertainty, at their most extreme. The three columns, from left to right, correspond to the residual wavefront errors for the ITM surface, ITM substrate, and end test mass (ETM) surface. The two rows, from top to bottom, correspond to the two scenarios with and without front-surface thermal imaging. For each panel, the residual RMS wavefront error of the mirror map is calculated, weighted by the Gaussian intensity profile of the main laser beam. The residual RMS wavefront error is smaller by a factor of 5-10 in the maps with thermal-imaging-based sensing, compared to the corresponding cases without it, demonstrating a significant improvement in the capability to correct thermally-driven wavefront distortions in both reflection and transmission.
4 Astrophysical Impact
[…]
Given a true set of wavefront errors due to laser beam heating, we use Monte Carlo to capture our practical inability to perfectly resolve those wavefront errors, which thus limits the precision to which we can identify the optimal power settings of the wavefront actuators and the optimal beam positioning. In each trial, we randomly sample the laser beam position, the ring heater power setting, and the FROSTI power setting for each of the four test masses. These parameters are drawn from independent Gaussian distributions centered on their optimal values, corresponding to perfectly centered beams and the wavefront actuator powers which minimize the full-aperture RMS wavefront errors. The standard deviations of the parameters, which are estimated in §3, represent the level of inferential uncertainty in practically determining each parameter’s optimal value.
@@ 4 Astrophysical Impact
Figure 4 shows the projected impact of the new wavefront sensing capabilities on the key performance metrics that determine LIGO’s quantum-limited sensitivity. The two cases, with (orange) and without (blue) thermal imaging of the test masses, are each obtained from roughly 1000 Monte Carlo trials assuming the inferential uncertainties described above. We find that 1000 trials are adequate to fully sample the distributions, based on convergence testing of their 95th percentiles. The left panel shows the probability distributions of laser power in the detector’s arm cavities for a fixed input power of 125 W. The middle panel shows the probability distributions of observed squeezing at 1 kHz for a fixed injected squeezing level of 9 dB, an effective level chosen to account for all attenuation losses and phase noise from the A+ squeezer system. The right panel shows the probability distributions of total resulting strain noise at 1 kHz.
As is visible in the left two panels of figure 4, there exist some cases in which trade-offs between optimizing the arm power (via the power recycling cavity coupling) and squeezing (via the signal recycling cavity coupling) lead to slightly better performance in one of these metrics, individually, without thermal imaging. However, the overall quantum noise level is consistently reduced in all cases when thermal imaging is employed, as shown in the right panel. We find that the precision inference enabled by thermal imaging of the test masses dramatically narrows the range of possible sensitivity outcomes, significantly improving the worst possible outcome at 95% confidence.
Figure 5: Comparison of the worst-case strain sensitivities of a LIGO A+ detector, at 95% confidence, with and without thermal-imaging-based inference. These two cases correspond to the two cases denoted by the dashed vertical lines in figure 4.
This large difference in worst-case scenarios translates directly into a statistical improvement in the expected strain sensitivity of the detector, which is shown in figure 5 for the worst observed outcomes in 95% of trials. That is, with 95% confidence, the detector will achieve at least the strain sensitivity indicated by these curves. Each case assumes the nominal A+ input power of 125 W and an effective injected level of 9 dB of frequency-dependent squeezing, as well as the nominal A+ thermal and technical noise contributions which are combined with the quantum noise curves calculated using Finesse. We find that thermal imaging of the test masses can enable up to a 34% greater reduction of the noise floor in LIGO A+, with maximum impact at the highest frequencies (which are quantum shot noise limited), corresponding to an increase of 11 Mpc in the sky-averaged detection range for binary neutron star mergers.
Finally, figure 6 shows the LIGO detector’s projected performance as the laser power is raised beyond the nominal A+ arm power target of 750 kW, towards the higher power levels required for LIGO $\rm A^{\#}$ and Cosmic Explorer. Each panel contains two shaded regions representing the range of outcomes spanned by the best 95% of trials in each of the two cases, with (orange) and without (blue) thermal imaging of the test masses. The left panel compares the possible ranges of input power required to achieve a given arm power, at 95% confidence. The middle and right panels compare the ranges of possible squeezing and strain sensitivity outcomes, for a given arm power level, at 95% confidence. Without the new inferential capabilities provided by thermal imaging, the probability of significant squeezing and power degradations grows rapidly above 750 kW, which is exacerbated by higher-order spatial modes becoming co-resonant in the recycling cavities with the changing thermal state of the ITMs. With the new wavefront sensing capabilities, the quantum noise performance of the detector is limited only by the intrinsic capability of the wavefront actuators to match the spatial profiles of the distortions.
5 Discussion
Our analysis shows that real-time precision inference of the thermal state of the optics is vital to achieving greater quantum-limited sensitivity in gravitational-wave detectors. We have demonstrated that front-surface thermal imaging offers a practical solution. The requirements for full-aperture wavefront correction are most stringent for the ITMs, due to wavefront distortions shifting the higher-order mode resonances of the signal recycling cavity [20], and the sensing requirements identified in this paper can satisfy them. Our assumption of a thermal steady-state FEA model is a realistic assumption for future detectors with the introduction of the CHETA offline laser heating system (see §1). We anticipate that such an offline heating system will be extended to all four test masses by LIGO $\rm A^{\#}$ and will be incorporated into the baseline Cosmic Explorer design. The impact of this work will thus extend far beyond A+, as it also delivers a key enabling sensing capability for next-generation gravitational-wave observatories.
@@ 5 Discussion
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
[…]